# RL-100: Performant Robotic Manipulation with Real-World Reinforcement Learning

### v2

@@ 4.4 Real-world RL
Real-world RL trains directly on real robot dynamics, optimizing deployment metrics (reliability, speed, safety) and yielding robust performance that continually adapts to disturbances—without sim-to-real gaps. Critical requirements include sample efficiency, stability under high-dimensional perception, safe continuous operation, and automated reward/reset mechanisms. While early work demonstrated end-to-end visuomotor learning (Levine et al., 2016) and large-scale grasping with off-policy methods (Kalashnikov et al., 2018), subsequent advances established key algorithmic foundations: off-policy actor-critic methods (SAC, TD3) for data efficiency (Haarnoja et al., 2018; Fujimoto et al., 2018), model-based approaches for sample acceleration (Chua et al., 2018; Janner et al., 2019), reset-free learning for autonomous operation (Eysenbach et al., 2018; Gupta et al., 2021), and learned reward specifications from visual classifiers or human feedback (Singh et al., 2019; Christiano et al., 2017). Despite these advances, most systems required extensive engineering, task-specific tuning, or long training times to achieve reliable performance.
@@
These scattered advances converge in SERL (Luo et al., 2024), a comprehensive framework that integrates high update-to-data ratio off-policy learning, automated resets, and visual reward specification to achieve several manipulation tasks. However, SERL relies solely on demonstrations and struggles with tasks requiring precision or recovery from failures. HIL-SERL (Luo et al., 2025) addresses these limitations by incorporating real-time human corrections during training, enabling the policy to learn from mistakes and achieve perfect success rates across diverse tasks, including dual-arm coordination and dynamic manipulation.
@@
While SERL and HIL-SERL report impressive on-robot learning efficiency and reliability on well-scoped tabletop skills, their evaluations typically employ action–space shaping (e.g., limiting wrist rotations and encouraging near-planar end-effector motion) and focus on short-horizon regimes with relatively low-dimensional control (Luo et al., 2024, 2025). Such constraints are pragmatic for safety and sample efficiency, but they reduce policy expressivity and can cap performance on orientation-critical, contact-rich, or compositionally complex tasks. In everyday home and factory scenarios, many skills inherently require full SE(3) control and substantial reorientation, including: deformable manipulation with twist and regrasp (e.g., towel folding), insertion/ejection in confined cavities with large tilt changes (e.g., orange juicing), fluid and granular control that hinges on container tilt (e.g., controlled pouring), dynamic release and trajectory shaping (e.g., agile bowling), cable routing or cloth placement with out-of-plane rotations, and bimanual reorientation. [^1] We do not claim these systems cannot be extended to such settings; rather, the reported experiments emphasize rapid learning on well-scoped tasks, leaving broader generalization, long-horizon composition, and orientation-sensitive control less explored. In contrast, our system retains full 6-DoF control without hard rotation constraints and targets these under-explored regimes by (i) using a diffusion/consistency visuomotor policy to capture diverse human strategies, (ii) unifying offline-to-online improvement with an OPE-gated PPO-style objective for nearly-monotonic improvement, and (iii) enabling high-frequency control via one-step consistency distillation, across dual-arm, deformable, and dynamic tasks with larger cross-object generalization.
@@ 5.1 Overview
Key specifications for observation and action spaces of all tasks are summarized in Tab. 2. We randomize initial object placements for each trial to encourage policy generalization, with specific ranges for each task depicted in Fig. 5. Two types of reward structure are used for our tasks. For all tasks except Dynamic Push-T, we use a sparse reward function where the agent receives a reward of +1 upon successful task completion and 0 at all other timesteps, labeled by a human supervisor pressing a keyboard. The Dynamic Push-T task, which requires continuous, high-precision control, utilizes a dense shaped reward. The total reward at each timestep $t$ is $r_{t}=r_{\text{pose}}+r_{\text{static}}+r_{\text{smooth}}$, where $r_{\text{pose}}=\exp(-3e)-1$ punishes the SE(3) discrepancy $e$ between the T-block and the desired pose; $r_{\text{static}}=-1$ if the movement of T-block at one timestep is below a given threshold and otherwise 0; $r_{\text{smooth}}=-5\|a_{t}-a_{t-1}\|^{2}_{2}$ punishes jerky actions. More details about our setup (calibration, point-cloud processing, low-level control) are provided in supplementary materials.
@@ 7 Simulation Experiments
To complement our real-robot evaluations, we benchmark RL-100 in simulation and compare it against state-of-the-art diffusion/flow–based offline-to-online RL methods, including DPPO (Ren et al., 2024), ReinFlow (Zhang et al., 2025b), and DSRL (Wagenmaker et al., 2025). We also conduct ablations to verify the contribution of each component in RL-100.
@@ 7.1 Main Results
We evaluate on three standard continuous-control suites that cover both dexterous manipulation and locomotion: Adroit (Rajeswaran et al., 2018), Meta-World (Yu et al., 2020), and MuJoCo locomotion (Fu et al., 2020; Todorov et al., 2012). Unless noted otherwise, observations include robot proprioception together with visual inputs (RGB frames or point clouds), and actions are continuous joint/EE commands. We report normalized returns and task success rates (mean $\pm$ std over multiple seeds) and compare RL-100 with baselines finetuning diffusion/flow policies as well as their offline/online training protocols.
@@ 8 Conclusion and Future Work
This paper focuses on the core challenges of real-world deployment of robot learning, which demands systems with high reliability, efficiency, and robustness. To this end, we present RL-100, a three-stage real-world RL training framework built atop a diffusion-based visuomotor policy, designed to “start from human priors” and then “go beyond human performance” in robotic manipulation. By unifying imitation learning and reinforcement learning under a single policy optimization objective, RL-100 preserves the strengths of human teleoperation demonstrations while continuously self-improving through autonomous exploration. In particular, a shared PPO-style policy-gradient loss is applied across both offline and online RL phases to fine-tune the diffusion policy’s denoising process, providing stable, near-monotonic performance gains. Moreover, the framework introduces a consistency distillation technique that compresses the multi-step diffusion policy (requiring K=5K=5–10 denoising steps) into a single-step consistency model, enabling high-frequency control at deployment without sacrificing performance. This unification of IL and RL under one objective, coupled with the one-step distilled policy, yields a controller that is both efficient to fine-tune and fast to execute in real-world settings.
RL-100 ’s efficacy was demonstrated across a diverse suite of real-world tasks, encompassing dynamic object pushing, dual-arm deformable cloth folding, agile high-speed bowling, precision pouring of granular and fluid materials, unscrewing, and a multi-stage orange juicing task. Deployment-level results show that the fully-trained RL-100 policy achieved 100% success rates on every task, attaining virtually perfect reliability in extensive evaluations. Notably, RL-100 also improved task efficiency, often matching or even surpassing skilled human operators in time-to-completion on multiple tasks. The trained policies proved robust over long durations, maintaining stable performance over 2+ hours of continuous operation, which attests to the framework’s long-horizon reliability. These results collectively indicate that RL-100 satisfies the stringent reliability, efficiency, and robustness requirements needed for real-world deployment.
Beyond excelling under nominal conditions, RL-100 exhibited strong generalization and robustness in the face of new variations and disturbances. Without any retraining, a single policy maintained high performance under significant environmental shifts. Across several such unseen scenario variations, RL-100 attained an average 90.0% success rate without any fine-tuning, underscoring the policy’s robustness to modest domain shifts. Furthermore, the framework’s consistency-model variant (the one-step distilled policy) demonstrated that high-frequency control is possible without loss of accuracy, as it too reached a 100% success rate on all evaluated tasks, including 250/250 successes in the challenging dual-arm towel folding. This blend of reliability and adaptability indicates that RL-100 can handle the kind of dynamic perturbations and variability inherent in real home or factory environments, rather than overfitting to narrow training conditions.
In summary, RL-100 integrates imitation learning and reinforcement learning into a unified paradigm that yields deployment-grade performance on real robots. It leverages human demonstration data as a foundation and then aligns to human-grounded objectives while exceeding human-level reliability through iterative self-improvement. The framework’s near-perfect success rates, broad task generality, and sustained operational robustness represent a significant step toward deploying autonomous manipulators in unstructured environments. These outcomes “offer a credible path to deployment in homes and factories”, where robots must perform daily manipulation tasks with super-human consistency and safety.
Future Work. There are several promising directions to extend and enhance this work. A priority is to extend evaluation to more complex, cluttered, and partially observable scenes that better mirror the variability of homes and factories. This includes dynamic multi-object settings, occlusions, specular/transparent materials, changing illumination, and non-stationary layouts. Stress-testing the policy beyond the current benchmark—where we already observe near-perfect success, strong long-horizon stability, and human teleoperation-level time-to-completion on multiple tasks—will help quantify limits and failure modes under realistic deployment conditions.
Our results indicate that small diffusion policies can attain high reliability and efficiency after modest fine-tuning. Building on this, we plan to scale post-training to larger, multi-task, multi-robot Vision-Language-Action (VLA) models. Key directions include: (i) scaling laws for data/model size vs. real-robot sample efficiency; (ii) cross-embodiment and cross-task transfer in a single policy; and (iii) aligning large VLA priors with RL-100 ’s unified objective to retain zero-shot semantic generalization while preserving the high success rates we report.
Although our pipeline already incorporates conservative operation and stable fine-tuning, reset and recovery remain practical bottlenecks. We will investigate autonomous reset mechanisms—learned reset policies, scripted recovery behaviors, task-aware environment fixtures, and failure-aware action chunking—to minimize on-robot human effort during training and evaluation. A principled reset strategy should reduce downtime, smooth data collection, and further stabilize online improvement, complementing RL-100 ’s OPE-gated iterative updates.

### v3

@@ 4.4 Real-world RL
Real-world RL trains directly on real robot dynamics, optimizing deployment metrics (reliability, speed, safety) and yielding robust performance that continually adapts to disturbances—without sim-to-real gaps. Critical requirements include sample efficiency, stability under high-dimensional perception, safe continuous operation, and automated reward/reset mechanisms. While early work demonstrated end-to-end visuomotor learning (Levine2016EndToEnd) and large-scale grasping with off-policy methods (Kalashnikov2018QTOpt), subsequent advances established key algorithmic foundations: off-policy actor-critic methods (SAC, TD3) for data efficiency (Haarnoja2018SAC; Fujimoto2018TD3), model-based approaches for sample acceleration (Chua2018PETS; Janner2019MBPO), reset-free learning for autonomous operation (Eysenbach2018Leave; Gupta2021Reset), and learned reward specifications from visual classifiers or human feedback (Singh2019End; Christiano2017Preferences). Despite these advances, most systems required extensive engineering, task-specific tuning, or long training times to achieve reliable performance.
@@
These scattered advances converge in SERL (luo2024serl), a comprehensive framework that integrates high update-to-data ratio off-policy learning, automated resets, and visual reward specification to achieve several manipulation tasks. However, SERL relies solely on demonstrations and struggles with tasks requiring precision or recovery from failures. HIL-SERL (luosr) addresses these limitations by incorporating real-time human corrections during training, enabling the policy to learn from mistakes and achieve perfect success rates across diverse tasks, including dual-arm coordination and dynamic manipulation.
@@
While SERL and HIL-SERL report impressive on-robot learning efficiency and reliability on well-scoped tabletop skills, their evaluations typically employ action–space shaping (e.g., limiting wrist rotations and encouraging near-planar end-effector motion) and focus on short-horizon regimes with relatively low-dimensional control (luo2024serl; luosr). Such constraints are pragmatic for safety and sample efficiency, but they reduce policy expressivity and can cap performance on orientation-critical, contact-rich, or compositionally complex tasks. In everyday home and factory scenarios, many skills inherently require full SE(3) control and substantial reorientation, including: deformable manipulation with twist and regrasp (e.g., towel folding), insertion/ejection in confined cavities with large tilt changes (e.g., orange juicing), fluid and granular control that hinges on container tilt (e.g., controlled pouring), dynamic release and trajectory shaping (e.g., agile bowling), cable routing or cloth placement with out-of-plane rotations, and bimanual reorientation. [^1] We do not claim these systems cannot be extended to such settings; rather, the reported experiments emphasize rapid learning on well-scoped tasks, leaving broader generalization, long-horizon composition, and orientation-sensitive control less explored. In contrast, our system retains full 6-DoF control without hard rotation constraints and targets these under-explored regimes by (i) using a diffusion/consistency visuomotor policy to capture diverse human strategies, (ii) unifying offline-to-online improvement with an OPE-gated PPO-style objective for nearly-monotonic improvement, and (iii) enabling high-frequency control via one-step consistency distillation, across dual-arm, deformable, and dynamic tasks with larger cross-object generalization.
@@ 5.1 Overview
Key specifications for observation and action spaces of all tasks are summarized in Tab. 2. We randomize initial object placements for each trial to encourage policy generalization, with specific ranges for each task depicted in Fig. 5. Two types of reward structure are used for our tasks. For all tasks except Dynamic Push-T, we use a sparse reward function where the agent receives a reward of +1 upon successful task completion and 0 at all other timesteps, labeled by a human supervisor pressing a keyboard. To further optimize execution efficiency (i.e., reduce time-to-completion) and stabilize control, we add (i) a small per-timestep step penalty that discourages unnecessarily long trajectories, and (ii) a penalty on action jitter, implemented as the deviation between the current action and a moving average of recent actions over a fixed temporal window, which suppresses high-frequency fluctuations in the control signal. Moreover, The Dynamic Push-T task, which requires continuous, high-precision control, utilizes a dense shaped reward. The total reward at each timestep $t$ is $r_{t}=r_{\text{pose}}+r_{\text{static}}+r_{\text{smooth}}$, where $r_{\text{pose}}=\exp(-3e)-1$ punishes the SE(3) discrepancy $e$ between the T-block and the desired pose; $r_{\text{static}}=-1$ if the movement of T-block at one timestep is below a given threshold and otherwise 0; $r_{\text{smooth}}=-5\|a_{t}-a_{t-1}\|^{2}_{2}$ punishes jerky actions. More details about our setup (calibration, point-cloud processing, low-level control) are provided in supplementary materials.
@@ 7 Simulation Experiments
To complement our real-robot evaluations, we benchmark RL-100 in simulation and compare it against state-of-the-art diffusion/flow–based offline-to-online RL methods, including DPPO (ren2024diffusion), ReinFlow (zhang2025reinflow), and DSRL (wagenmaker2025steering). We also conduct ablations to verify the contribution of each component in RL-100.
@@ 7.1 Main Results
We evaluate on three standard continuous-control suites that cover both dexterous manipulation and locomotion: Adroit (adroit), Meta-World (metaworld), and MuJoCo locomotion (d4rl; mujoco). Unless noted otherwise, observations include robot proprioception together with visual inputs (RGB frames or point clouds), and actions are continuous joint/EE commands. We report normalized returns and task success rates (mean $\pm$ std over multiple seeds) and compare RL-100 with baselines finetuning diffusion/flow policies as well as their offline/online training protocols.
@@ 8 Conclusion and Future Work
The results advance the state of the art in robotic manipulation in terms of reliability, efficiency, and robustness. Across seven real-world tasks—dynamic pushing, dual-arm soft-towel folding, agile bowling, pouring (granular/fluids), unscrewing, and multi-stage juicing—it achieves 100% success in 900/900 trials. The policy generalizes across initial object placements and tolerates shape/size variations, while matching or surpassing expert teleoperators in time-to-completion. It remains stable in extended deployment, including a public mall demo running more than seven hours without failure, and sustains high performance under significant environmental shifts without retraining. Together, these results indicate that RL-100 meets deployment-level requirements. In summary, RL-100 uses a small set of human demonstrations as a foundation and then aligns with human-grounded objectives, exceeding human-level reliability through iterative self-improvement. This reduces the burden on data-collection staff and further benefits from large volumes of autonomous on-robot rollouts. Its perfect success rates, broad task generality, and sustained operational robustness represent a significant step toward deploying autonomous manipulators in unstructured environments, outlining a credible path to deployment in homes and factories. The findings suggest that small diffusion policies can attain high reliability and efficiency after modest fine-tuning. Because RL-100’s training pipeline is model-agnostic, we can scale post-training to larger, multi-task, multi-robot diffusion/flow-based vision-language-action (VLA) models. Key directions include: (i) establishing scaling laws for data and model size versus real-robot sample efficiency; (ii) cross-embodiment and cross-task transfer within a single policy; and (iii) aligning large VLA priors with RL-100’s unified objective to retain zero-shot semantic generalization while preserving the reported high success rates. We will also extend evaluation to more complex, cluttered, and partially observable settings that better mirror homes and factories, including dynamic multi-object scenes, occlusions, specular/transparent materials, changing illumination, and non-stationary layouts. Although our pipeline already emphasizes conservative operation and stable fine-tuning, reset and recovery remain practical bottlenecks. We will explore autonomous reset mechanisms—learned reset policies, scripted recovery behaviors, task-aware fixtures, and failure-aware action chunking—to minimize on-robot human effort during training and evaluation.
@@ Acknowledgement
9 Appendix
9.1 Experiment Setting
This section presents the setup and methodology for conducting real-world robot manipulation experiments. The experimental framework is designed to integrate perception, control, and demonstration collection in a coherent pipeline. We employ the Intel RealSense L515 camera to capture depth images, which are subsequently processed into 3D point clouds and transformed into the robot’s root frame for consistent spatial representation. Intrinsic and extrinsic calibrations ensure accurate mapping between image coordinates and the robot workspace, while spatial filtering and downsampling produce compact yet informative point cloud observations suitable for high-frequency control.
The robotic platform consists of three manipulators—UR5, xArm, and Franka Emika Panda—equipped with dexterous or gripper-type end-effectors, as well as a passive fixed effector. Asynchronous control schemes are implemented across all manipulators to achieve low-latency, high-frequency execution, with task-specific interpolation strategies ensuring smooth and responsive motion.
Demonstrations for seven manipulation tasks are collected using teleoperation interfaces matched to task complexity. For dexterous 3D tasks, the Apple Vision Pro provides natural hand tracking, which is converted into robot joint or Cartesian commands via inverse kinematics for the LeapHand. For planar or low-dimensional tasks, a standard game joystick is used to provide delta-action control of the end-effector in the plane. Each task includes approximately 100 demonstration trajectories, efficiently recorded within a few hours per task.
Overall, the experimental setup provides a consistent, flexible, and practical framework for acquiring high-quality real-world data and evaluating robot manipulation policies across a diverse set of tasks.
9.1.1 Camera Calibration
In the real-world experiments, we employ the Intel RealSense L515 camera, chosen for its ability to provide high-fidelity 3D point cloud observations owing to its superior depth accuracy. This ensures reliable perception for robot manipulation tasks across diverse environments.
The intrinsic calibration is performed using a Charuco board, which serves as the reference pattern for estimating the camera’s internal parameters. Multiple images of the board are captured from varying angles and distances. The board corners are then detected and used to compute the intrinsic matrix and distortion coefficients through an optimization process that minimizes the reprojection error. This procedure guarantees an accurate mapping between 2D image coordinates and their corresponding 3D spatial points.
The extrinsic calibration aligns the camera frame with the robot’s root frame using an AprilTag placed on the tabletop. First, a manual measurement establishes the transformation $T_{\text{tag2root}}$, which encodes the pose of the tag relative to the robot base. Subsequently, the AprilTag is detected in the camera’s RGB channel, yielding the transformation $T_{\text{tag2camera}}$ from the tag to the camera frame. The camera-to-root transformation is then obtained as:
$T_{\text{camera2root}}=T_{\text{tag2root}}\cdot(T_{\text{tag2camera}})^{-1}$
This calibrated extrinsic transformation is stored and utilized throughout the experiments to maintain consistent spatial alignment between the camera observations and the robot workspace.
9.1.2 Point Cloud Processing
The raw depth maps captured by the RealSense L515 are first back-projected into 3D point clouds using the calibrated intrinsic parameters, thereby establishing a metric reconstruction of the scene in the camera frame. These point clouds are then transformed into the robot’s root frame using the calibrated extrinsic transformation $T_{\text{camera2root}}$, ensuring spatial consistency between visual observations and the robot’s operational workspace.
To suppress environmental noise and spurious reflections, a spatial bounding box is applied in the root frame, filtering out irrelevant regions while preserving only the volume above the tabletop. This step enhances the semantic relevance of the retained points, focusing on task-relevant structures.
Finally, to improve computational efficiency while maintaining geometric fidelity, farthest point sampling (FPS) is applied. The processed point cloud is uniformly downsampled to 512 points, yielding a compact yet informative representation suitable for robot perception and policy learning. During deployment, the camera operates in a dedicated thread with a buffered pipeline, enabling asynchronous acquisition of observations. This design minimizes latency and allows the effective point cloud update rate to exceed 25 Hz, ensuring timely and reliable perception for closed-loop control.
9.1.3 Robotic Arm & End-effector Control
The experimental platform incorporates three robotic manipulators—UR5, xArm, and Franka Emika Panda—equipped with different types of end-effectors, including the LeapHand dexterous hand, the Robotiq 2F-85 gripper, and a custom 3D-printed fixed effector. All control modules are designed to operate asynchronously, ensuring low-latency communication and high-frequency execution for responsive and stable manipulation.
For the UR5, relative action commands are transmitted at 30 Hz. These commands are handled through the Real-Time Data Exchange (RTDE) interface, where they are interpolated to 125 Hz for execution. A lookahead time of $1/30$ s and a control gain of 200 are used, providing a balance between smooth motion trajectories and timely responsiveness.
For the xArm, control is performed via absolute Cartesian position commands at 30 Hz. Linear interpolation is applied to the position trajectories, while orientation is represented in quaternion form and interpolated via spherical linear interpolation (slerp). The resulting trajectory is executed at 200 Hz, yielding continuous and smooth motion in task space.
For the Franka Emika Panda, absolute Cartesian position commands are also issued at 30 Hz. The robot’s low-level controller performs trajectory interpolation, converting the input into control signals at 1000 Hz, thereby ensuring precise and compliant execution.
Regarding the end-effector control, both the LeapHand and the Robotiq 2F-85 are driven asynchronously at 30 Hz. The policy performs end-to-end joint control for the LeapHand, while the gripper uses binary control for the Soft-towel Folding task and continuous control for the Orange Juicing task. The custom 3D-printed fixed effector, by contrast, is purely passive and requires no active control, serving as a simple but reliable interface for specific experimental settings like Dynamic Push-T or Agile Bowling.
9.2 Data Collection
Data for seven real-world manipulation tasks are collected using teleoperation interfaces selected to match the complexity and dimensionality of each task. These tasks include both dexterous three-dimensional manipulations—such as folding, pouring, unscrewing, and juicing—and planar or constrained motions, such as Dynamic Push-T and Agile Bowling. The data collection strategy aims to provide sufficient coverage of the task space while maintaining practical efficiency in demonstration acquisition.
For tasks involving complex 3D hand motions, the Apple Vision Pro is employed. Its built-in spatial hand-tracking captures natural hand and finger movements, which are mapped to robot end-effector motions. Before starting each demonstration, the operator’s hand pose is aligned with the robot end-effector through an initialization procedure, ensuring a consistent reference frame. During teleoperation, wrist and finger motions are streamed in real time and converted into robot joint or Cartesian commands, while gripper grasping actions are triggered via pinch gestures. For the LeapHand dexterous hand, inverse kinematics is solved in PyBullet to produce feasible joint configurations from the tracked hand poses. This setup allows operators to efficiently provide demonstrations for tasks requiring dexterous, spatially rich manipulations.
For planar or low-dimensional manipulation tasks—such as Dynamic Push-T and Agile Bowling—a standard game joystick is used. The joystick’s XY-axis values are mapped to delta actions of the robot end-effector in the plane. This approach provides an intuitive and efficient means for recording demonstrations where precise two-dimensional control is sufficient.
Our dataset is intentionally varied. Some tasks feature a large number of demonstrations to test performance limits, while others use smaller datasets to evaluate sample efficiency. Tab. 7 provides a complete summary.